OpenStudy (anonymous):

Solve Log(5)2x-5= -4

OpenStudy (mertsj):

What is the base?

OpenStudy (anonymous):

5

OpenStudy (mertsj):

What is the exponent?

OpenStudy (anonymous):

-4 I believe

OpenStudy (mertsj):

Then write 5^-4=2x-5 and solve.

OpenStudy (anonymous):

Is it 2.50 when rounded to the nearest hundredth?

OpenStudy (mertsj):

\[5^{^{-4}}=2x-5\] \[\frac{1}{625}=2x-5\] \[5 \frac{1}{625}=2x=\frac{3126}{625}\]

OpenStudy (mertsj):

\[x=\frac{1563}{625}\]

OpenStudy (mertsj):

x=2.5008

OpenStudy (anonymous):

Thank you! =)

OpenStudy (mertsj):

yw